An air-conditioning system operating on the reversed Carnot cycle is



required to transfer heat from a house at a rate of 750 kJ/min to



maintain its temperature at 24°C. If the outdoor air temperature is



35°C, determine the power required to operate this air-conditioning



system.

1
Expert's answer
2022-02-28T11:09:24-0500

Solution;

Given;

"Q=750kJ\/min"

"T_s=35\u00b0c"

"T_h=25\u00b0c"

"C.O.P=\frac{T_h}{T_s-T_h}=\frac{25+273}{35-25}=29.8"

Also;

"C.O.P=\frac{Q}{W}"

"W=\frac{Q}{C.O.P}=\frac{750}{29.8}=25.168kJ/min"